how many pounds of cashews costing $8.00 per pound must be mixed with peanuts costing $3.00 per pound to make a 25 pound mixture that costs $5.00 per pound
Answer by ikleyn(53763)

.
Let x be the amount (in pounds) of cashews costing $8.00 per pound to be mixed.

Then the amount of peanuts costing $3.00 per pound is (25-x)

The equation based on the price of the mixture is

8x + 3(25-x) = 25*5.


Simplify and solve for x:

8x + 75 - 3x = 125,

5x = 50   --->   x = {50/5}}} = 10.


Answer.  10 pounds 0f the $8.00 cashew is to be mixed with 15 pounds of the $3.00 peanuts.


Chheck.  $8*10 + $3*15 = 80 + 45 = 125 dollars.

          $25*5 = 125 dollars.  Correct !!


There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for dry substances like coffee beans, nuts, cashew and peanuts
    - Word problems on mixtures for dry substances like candies, dried fruits
    - Word problems on mixtures for dry substances like soil and sand
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
    - OVERVIEW of lessons on word problems for mixtures
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Mixture problems